OpenStudy (flowers):

Can someone help me with these 3x3 Linear Systems 5x+2y+z=4 X+2z=4 2x+y-z=-1 3x+10y-12z X-5y=0 X-4z=0

OpenStudy (tkhunny):

What is your favorite method of solution?

OpenStudy (flowers):

Elimination

OpenStudy (tkhunny):

Okay, what's preventing you from doing that? x + 2z = 4 looks like a good place to start. 1) Multiply it by -5 and add to the first equation. 2) Multiply it by -2 and add it to the third equation. You will have eliminated x from all but equation #2.

OpenStudy (flowers):

Teacher say try to use substitutions but i really dont like to use substitutions I didnt realize that i can use elimination

OpenStudy (tkhunny):

Yes, elimination is just fine. I like to say, "Unique Answers don't care how you find them."

563blackghost (563blackghost):

I can help but I use substitution best.

563blackghost (563blackghost):

Is that alright @flowers ?

OpenStudy (tkhunny):

It should be pointed out that the first equation int eh second set is incomplete.

OpenStudy (flowers):

Sure I need to learn it any way :)

563blackghost (563blackghost):

I haven't quite taken a look at the second set only the first one.

563blackghost (563blackghost):

Alrighty :) Due to the fact that the second equation `x+2z = 4` contains two variables we would change this into a function of x. To do this we would subtract 2z from each side. \(\LARGE\bf{x+2z-2z=-2z+4 \rightarrow x = -2z + 4}\)

563blackghost (563blackghost):

Now we would input this equation of x into the two other equations. We would simplify them. \(\huge\bf{5(-2z+4)+2y+z=4}\) What would this simplify to @flowers ?

563blackghost (563blackghost):

Remember to first distribute into the parenthesis and then combine like terms.

OpenStudy (flowers):

Sorry open study likes to reshape it self 10 +20

563blackghost (563blackghost):

\(\LARGE\bf{-10z+20+2y+z=4}\) Combine like terms. \(\LARGE\bf{-9z + 2y + 20 = 4}\) Now we need the numbers without variables to be on one side. So we have to subtract 20 from both sides. \(\LARGE\bf{-9z + 2y = 20 \color{red}{-20} = 4 \color{red}{-20}}\) \(\huge\bf{-9z + 2y = -16}\) Do you understand?

563blackghost (563blackghost):

@flowers

563blackghost (563blackghost):

So we leave this equation alone (we will use later). Now we input the equation of x into the 3rd equation. \(\huge\bf{ 2(-2z+4)+y-z=-1}\) What would this simplify to? @flowers

563blackghost (563blackghost):

So sorry that equal sign near 20 is wrong its to be + not =. Darn the shift on my keyboard xD

OpenStudy (flowers):

4z+8

563blackghost (563blackghost):

Nice :) Now we combine like terms. \(\LARGE\bf{-5z+y+8=-1}\) Now the ones without a variable go to one side so we subtract 8 from both sides. \(\LARGE\bf{-5z + y + 8 \color{red}{-8} = -1 \color{red}{-8}}\) \(\LARGE\bf{-5z+y=-9}\)

563blackghost (563blackghost):

Now we have. \(\huge\bf{-9z+2y=-16}\) and \(\huge\bf{-5z + y = -9}\) We will change the second equation into an equation of y since it is to itself. So we add 5z to each side. \(\huge\bf{y = 5z -9}\) ~~~~~~~~~~~~~~~ Now we would input this equation of y into the first equation that we simplified. \(\huge\bf{-9z + 2(\color{red}{5z-9})=-16}\) What does that simplify to?

OpenStudy (flowers):

7z(-4z)=16

OpenStudy (flowers):

I think I'm way off

563blackghost (563blackghost):

? How did you come by 7z(-4z) = 16? Remember when distributing you multiply with each term in the parenthesis. \(\huge\bf{5z \times 2 -9 \times 2 = ?}\)

OpenStudy (flowers):

Dang it Looking at the wrong thing

OpenStudy (flowers):

Funny I have the right idea on my paper but not here

563blackghost (563blackghost):

Its fine :) Just make sure to keep your mind focus otherwise it will go off and problems will occur.

563blackghost (563blackghost):

So we have. \(\huge\bf{-9z + (5z \times 2 - 9 \times 2)=-16}\)

OpenStudy (flowers):

Woah Multiple 5 and 2 right?

563blackghost (563blackghost):

Yes :) First do the parenthesis.

563blackghost (563blackghost):

But to be accurate `multiply 5z and 2`

OpenStudy (flowers):

OK -9z+10z-18=-16

563blackghost (563blackghost):

Nice :) Now combine like terms. \(\huge\bf{-9z + 10z - 18 = -16}\) The like terms are `-9z and 10z`.

OpenStudy (flowers):

18z-18=-16

OpenStudy (flowers):

I mean 19z

OpenStudy (flowers):

Wait its a negative number Ugh

563blackghost (563blackghost):

Not quite. \(\huge\bf{-9z + 10z = z}\) Since we add -9 to a positive being 10 it will lower the number to 1. So we have. \(\huge\bf{z-18=-16}\) Now we need z to itself so we add 18 to each side since it is being added to z. \(\huge\bf{z-18 \color{red}{+18}=-16 \color{red}{+18}}\)

563blackghost (563blackghost):

What does z equal?

OpenStudy (flowers):

Z=2

563blackghost (563blackghost):

Yes :)

OpenStudy (flowers):

Seriously wow omg thank you so much

563blackghost (563blackghost):

Now we need to input z into the equation that we turned into an equation of y `y = 5z -9`. \(\huge\bf{y=5(2)-9}\) What does y equal?

563blackghost (563blackghost):

Not done yet ^.^ still need to find x and y.

OpenStudy (flowers):

3-9

563blackghost (563blackghost):

Not quite. \(\huge\bf{y=10-9}\)

OpenStudy (flowers):

Ik Oops thinking differently Meant to say 10

563blackghost (563blackghost):

Thats alright ^.^ Though what is `10-9`?

OpenStudy (flowers):

1

563blackghost (563blackghost):

Correct so y = 1 :) Now we need to find x.

563blackghost (563blackghost):

Now we input y and z into the first equation `(5x + 2y + z = 4)`. \(\huge\bf{5x+2(1)+2=4}\) Simplify :)

OpenStudy (flowers):

Okay :) 5x+2+2=4

563blackghost (563blackghost):

Nice :) Now we combine like terms `being 2 + 2`. \(\huge\bf{5x+4=4}\) Now we need `5x` to itself so we would subtract 4 from both sides. \(\huge\bf{5x+4 \color{red}{-4} = 4 \color{red}{-4}}\) What does that equal?

OpenStudy (flowers):

5x=0 /5 /5 X=0

563blackghost (563blackghost):

Nice :) Due to the fact it is formatted as \(\large\bf{(x,y,z)}\) your answer would be as such... \(\huge\bf\color{red}{(0,1,2)}\)

OpenStudy (flowers):

Yay thank you so much I can now do the other problems ;)

563blackghost (563blackghost):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @tkhunny It should be pointed out that the first equation in the second set is incomplete. \(\color{#0cbb34}{\text{End of Quote}}\) I can not really help for the second set due to the fact the first is incomplete as hunny said. `"3x+10y-12z"` it seems to be missing its total.

OpenStudy (flowers):

Ikr but this algebra 2 so yea..

563blackghost (563blackghost):

Are you sure its only... `3x + 1y - 12z` in order to solve 3x3 systems you need to have a total in each equation. Are you sure its not missing something. `3x + 1y - 12z = ?`

OpenStudy (flowers):

Let me check

OpenStudy (flowers):

Oops 3x+10y-12z=40

563blackghost (563blackghost):

Ooo ok good to know. I know this is bad timing but I must take my leave for the afternoon. @tkhunny can you help @flowers ?

563blackghost (563blackghost):

or @Directrix @zepdrix @mathmate ?

OpenStudy (flowers):

Oh no Your fine I can do it

OpenStudy (flowers):

Thank you for your help

563blackghost (563blackghost):

Oh ok. Well good luck with your work and I am glad I was able to help ^.^ See ya around.

563blackghost (563blackghost):

Your very welcome :)